# Causal Inference with Large Language Model: A Survey

### v1

@@ 1.1 NLP, LLM, and Causality
Causal inference is an important area to uncover and leverage the causal relationships behind observations, enabling a deep understanding of the underlying mechanism and potential interventions in real-world data systems. Different from most classical statistical studies, causal inference presents unique challenges due to its focus on "causation instead of correlation", which intricates a complicated integration of human knowledge (e.g., domain expertise and common sense), mathematics, and data mining. Due to the inherent proximity to the human cognitive process, causal inference has become pivotal in many high-stakes domains such as healthcare Glass et al. (2013), finance Atanasov and Black (2016), and science Imbens and Rubin (2015).
Traditional causal inference frameworks, such as structural causal model (SCM) Pearl (2009) and potential outcome framework Imbens and Rubin (2015) have systematically defined causal concepts, quantities, and measures, followed up with multiple data-driven methods to discover the underlying causal relationships Spirtes and Zhang (2016); Nogueira et al. (2022); Vowels et al. (2022) and estimate the significance of causal effects Winship and Morgan (1999); Yao et al. (2021). Despite their success, there is still a large gap between existing causal methods and human’s judgment Kıcıman et al. (2023); Zečević et al. (2023); Jin et al. (2023a), covering different aspects such as the lack of domain knowledge, logic inference, and cultural background. Besides, most traditional causal inference approaches only focus on tabular data, lacking the ability to discover and utilize the causality inside natural language. However, the motivation for causal inference in natural language has persisted over an extended period, offering a multitude of potential applications. For example, clinical text data in electronic health records (EHR) contains a large amount of underlying causal knowledge that can be utilized for healthcare-related research. In general, causal inference in natural language processing (NLP) is a promising research path with strong motivation, which offers a spectrum of challenges and benefits concurrently. Recently, the burgeoning field of large language models (LLMs) has shed light on its potential to improve traditional causal inference, offering fresh perspectives to bridge the gap between human cognition and causal inference methodologies Feder et al. (2022).
@@ 1.2 Challenges of Causal Inference in NLP
Although LLMs have shown eye-catching success in various tasks, causal inference still presents many distinctive challenges for LLM capabilities. Different from regular data types, the nature of natural language brings difficulties in causal processing and analysis. As aforementioned, text data is often unstructured, high-dimensional, and large-scale, in which context traditional causal methods are not applicable. Besides, causal relationships inside the text are often obscure and sparse. The complicated semantic meaning and ambiguity hidden in text data require sophisticated language modeling technologies to discover clear causal relationships, and also entail hurdles for other causal tasks such as causal intervention and counterfactual reasoning. These challenges demand new perspectives, assumptions, and technologies to address them effectively, offering revolutionary opportunities for current causal inference studies.
1.3 Opportunites that LLMs Bring to Causal Inference
Despite the challenges, natural language has significant potential to yield advantages in causal inference. As LLMs have become increasingly sophisticated with diverse applications in recent years, the feasibility of understanding and unraveling causal relationships within linguistic data has been substantially improved. In general, LLM can bring many benefits to causal inference, including but not limited to the following main aspects:
Domain knowledge. Typical statistical methods for causal inference often only focus on the numerical values of variables, while in many scenarios, domain knowledge plays an important role in causality-related tasks as it provides us with additional information to discover the true causal relationships and make meaningful interventions. For example, in many scientific domains such as medicine, incorporating the domain knowledge can draw conclusions that cannot be obtained solely through pure statistical methods, and expedite the development of relevant fields. However, collecting domain knowledge from human experts often demands considerable effort. Fortunately, the recent developments in NLP and LLM can extract domain knowledge from large-scale text information and thereby facilitate causal inference.
Common sense. Similar to domain knowledge, language models can serve as an effective tool to learn and utilize humans’ general common sense to promote causal inference. As discussed in Kıcıman et al. (2023), a variety of common sense in different scenarios affects humans’ recognition of causal relationships. For example, logical reasoning is essential for causal inference in law cases. Besides, abnormal events are often more likely to be recognized as causes for an outcome of interest in common sense.
Sematical concept. Compared with regular data types, natural language contains nuances, variations, and the richness of human expression, requiring advanced techniques for semantic analysis. Therefore, grasping clear causal concepts and relationships from text data is much more challenging than other data types. Recent progress in NLP and LLM technologies, especially their ability in semantic modeling pave the way for in-depth causal studies in the next step.
Explainable causal inference. LLMs have the potential to offer sophisticated natural language-based tools that facilitate a more engaging and intuitive human understanding of causal inference. By leveraging these tools, users can gain clearer insights into the causal reasoning process and the resulting findings, making intricate concepts more accessible. This approach not only simplifies the interpretation of causal relationships but also enhances the ability of individuals to interact meaningfully with the causal inference results, thereby bridging the gap between complex analytical processes and user comprehension.
@@ 1.4 Contribution and Uniqueness
Contribution. This survey systematically reviews existing studies of using LLMs for causal inference. The main contribution of our survey can be summarized as:
  • We provide a well-structured categorization for existing studies in this area, organizing them into distinct groups based on their tasks (Section 2) and technologies (Section 3).
  • We show a detailed evaluation comparison of different existing LLMs (Section 4), and we elucidate the main observations, connections, and insights based on the results.
  • We present a comprehensive summarization of benchmark datasets in the field, covering many important aspects for further study, as illustrated in Table 1.
  • We discuss the limitations of current studies and future directions in Section 5. These discussions highlight gaps and further opportunities that have not been fully explored in existing literature, offering a novel perspective on potential advancements in the area.
Differences from existing surveys. There have been several related surveys of LLM applications in causal inference Liu et al. (2024b); Kıcıman et al. (2023). We summarize the differences between our survey and existing ones as follows: (1) Main scope. Our paper offers an extensive and comprehensive exploration of LLMs in causal inference, i.e., “LLMs for causality", different from some of existing surveys Liu et al. (2024b) that primarily focus on “causality for LLMs". (2) Structure and content. The structure of our survey is significantly different from previous surveys, including a well-organized presentation dedicated to tasks, methods, datasets, and evaluation, thereby offering a clearer and more thorough examination. (3) Up-to-date. Our survey incorporates cutting-edge research developments, providing an up-to-date snapshot of the latest progress and trends.
@@
4.2) [ Pairwise Causal Discovery [ GPT-4 (Jin et al.; Cai et al.; Liu et al.; Gao et al.; Zečević et al.;
@@ 2.1 Causality
Structural causal model. Structural causal model (SCM) Pearl (2009) is a widely used model to describe the causal relationships inside a system. An SCM is defined with a triple $(U,V,F)$: $U$ is a set of exogenous variables, whose causes are out of the system; $V$ is a set of endogenous variables, which are determined by variables in $U\cup V$; $F=\{f_{1}(\cdot),f_{2}(\cdot),...,f_{|V|}(\cdot)\}$ is a set of functions (a.k.a. structural equations). For each $V_{i}\in V$, $V_{i}=f_{i}(pa_{i},U_{i})$, where “$pa_{i}\subseteq V\setminus{V_{i}}$” and “$U_{i}\subseteq U$" are variables that directly cause $V_{i}$. Each SCM is associated with a causal graph, which is a directed acyclic graph (DAG). In the causal graph, each node stands for a variable, and each arrow represents a causal relationship.
Ladder of causation. The ladder of causation Pearl and Mackenzie (2018); Bareinboim et al. (2022) defines three rungs (Rung 1: Association; Rung 2: Intervention; Rung 3: Counterfactuals) to describe different levels of causation. Each higher rung indicates a more advanced level of causality. The first rung "Association" involves statistical dependencies, related to questions such as "What is the correlation between taking a medicine and a disease?". The second rung "Intervention" moves further to allow interventions on variables. Questions related to this rung are like "If I take a certain medicine, will my disease be cured?". The top rung "Counterfactuals" relates to imagination or retrospection queries like "What if I had acted differently?", "Why?". Answering such questions requires knowledge related to the corresponding SCM. Counterfactual ranks the highest because it subsumes the first two rungs. A model that can handle counterfactual queries can also handle associational and interventional queries.
@@ 2.2 Causal Tasks and Related Rungs in Ladder of Causation
Causal inference involves various tasks. Figure 1 shows an overview of causal inference tasks and their positions in the ladder of causation. We also show several examples of prompts corresponding to each rung. We list several main causal tasks which are most widely studied as follows:
@@
Causal effect estimation. Causal effect estimation (a.k.a. treatment effect estimation) targets on quantifying the strength of the causal influence of a particular intervention or treatment on an outcome of interest. In different scenarios, researchers may focus on the causal effect of different granularities, such as individual treatment effect (ITE, i.e., treatment effect on a specific individual), conditional average treatment effect (CATE, i.e., average treatment effect on a certain subgroup of population), average treatment effect on the treated group (ATT), and average treatment effect (ATE, i.e., average treatment effect on the entire population). Besides, people are also interested in the direct or indirect causal effects in certain scenarios, such as natural direct effect (NDE), controlled direct effect (CDE), and natural indirect effect (NIE). Another task related to causal effect estimation is backdoor adjustment (BAJ), which aims to block all backdoor paths Pearl (2009) from the treatment to the outcome to exclude non-causal associations. Causal effect estimation tasks often span over Rung 2 and Rung 3 in the ladder of causation.
Other tasks. There are many other tasks in causal inference. Among them, causal attribution (CA) refers to the process of attributing a certain outcome to certain events. Counterfactual reasoning (CR) investigates what might have happened if certain events or conditions had been different from what actually occurred. It explores hypothetical scenarios by considering alternative outcomes based on changes in “what if" circumstances. Causal explanation (CE) aims to generate explanations for an event, a prediction, or any causal reasoning process. This task often needs to answer causal questions in a specified human-understandable form or plain language. It is often in Rung 2 or Rung 3, depending on the specific context. It is worth mentioning that, in many cases, different causal tasks may exhibit natural overlap in their scope. For instance, causal attribution and explanation commonly intersect with causal discovery and causal effect estimation. However, each task maintains a distinct focus and emphasis.
@@
Table 1: Datasets for LLM-related causal inference, including publication year, applicable tasks (CD=causal discovery; Eff=causal effect estimation; CR=counterfactual reasoning; CE=causal explanation), dataset size (as different datasets are not in a consistent form, we show the size w.r.t. different units, where P=causal pairs; R=causal relations; S=samples; Q=questions), domain, generation process (R: real-world; S: synthetic), number of data sources, and citations.
@@ 3 Methodologies
Recently, there have emerged many efforts (Kıcıman et al., 2023; Chen et al., 2024a; Gao et al., 2023) to leverage LLMs for causal tasks. Different from traditional causal inference approaches which are either data-driven or based on expert knowledge, the nature of LLM training and adoption introduces novel methodologies in causal inference, offering new perspectives and insights for discovering and utilizing causal knowledge in future research and applications. A list of LLMs developed or evaluated in different causal tasks is shown in Figure 2. We summarize the current methodologies of LLMs for causal tasks into the following categories:
Prompting. Most existing works Chen et al. (2024a); Kıcıman et al. (2023); Long et al. (2023); Jin et al. (2023a) of causal reasoning with LLMs focus on prompting, as it is the most straightforward approach. This line of work includes both regular prompting strategies (such as basic prompt, In-Context Learning (ICL) Brown et al. (2020), and Chain-of-Thought (CoT) Wei et al. (2022)) and causality-specific strategies. For regular prompting, most studies directly use a basic prompt (i.e., directly describe the question without any example or instruction). There are also other efforts to devise more advanced prompting strategies. Among them, CaLM Chen et al. (2024a) has tested 9 prompting strategies including basic prompt, adversarial prompt Wallace et al. (2019); Perez and Ribeiro (2022), ICL, 0-shot CoT (e.g., “let’s think step by step” without any examples) Kojima et al. (2022), manual CoT (i.e., guide models with manually designed examples), and explicit function (EF) (i.e., using encouraging language in prompts) Chen et al. (2024a). Other works Kıcıman et al. (2023); Long et al. (2023); Gao et al. (2023); Ban et al. (2023) also design different prompt templates. These works show substantial improvement potential of prompt engineering in causal reasoning tasks. For example, results in Kıcıman et al. (2023); Chen et al. (2024a); Long et al. (2023) show adding simple sentences like "you are a helpful causal assistant" or "you are an expert in [DOMAIN NAME]" can impressively improve the causal inference performance for many models. Apart from these regular methods, other studies propose causality-specific prompting strategies. For example, CausalCoT Jin et al. (2023a) is a multi-step prompting strategy that combines CoT prompting and the causal inference engine Pearl and Mackenzie (2018).
@@
Combining LLMs with traditional causal methods. Another line of works combine LLMs with traditional causal methods. Considering causal inference often heavily relies on numerical reasoning, an exploration in Ban et al. (2023) leverages LLMs and data-driven causal algorithms such as MINOBSx Li and Beek (2018) and CaMML O’Donnell et al. (2006). This method outperforms both original LLMs and data-driven methods, indicating a promising future for combining the language understanding capability of LLMs and the numerical reasoning skills of data-driven methods in complicated causal tasks. Jiralerspong et al. Jiralerspong et al. (2024) combine LLM with a breadth-first search (BFS) approach for full causal graph discovery. It considers each causal relation query as a node expansion process, and gradually constructs the causal graph by traversing it using BFS. This method significantly reduces the time complexity from $O(n^{2})$ to $O(n)$, where $n$ is the number of variables. While it does not require access to observational data, their experiments show that the performance can be further enhanced with observational statistics.
Knowledge augmentation. LLMs with augmented knowledge can often better execute tasks for which they are not well-suited to perform by themselves. This idea is tempting for causal tasks as they often require corresponding causal knowledge. Pawlowski et al. (2023) propose two types of knowledge augmentation to answer causal questions. Their augmentations include a context augmentation that provides a causal graph and the ITEs in the prompt, and a tool augmentation with API access to an expert system that performs causal reasoning tasks. These augmented LLMs have promising results in answering causal queries. The tool augmentation performs more robustly over problems with different sizes, likely because LLM does not need to reason through the graph itself, but can instead utilize the API for causal reasoning. LACR Zhang et al. (2024) applies retrieval augmented generation (RAG) to enhance the knowledge base of LLM for causal discovery tasks, where the knowledge resources are from a large scientific corpus containing hidden insights about associational/causal relationships. Similarly, Samarajeewa et al. (2024) use causal graphs as external sources for causal reasoning. Noticeably, knowledge augmentation is particularly useful for causal tasks in professional domains. For example, RC2R Yu et al. (2024) combines LLMs with expert knowledge in financial knowledge graphs to analyze the causal mechanisms of financial risk contagion. CausalKGPT Zhou et al. (2024a) is a causal knowledge graph augmented LLM, designed to handle causal reasoning for quality defects in aerospace product manufacturing.

### v2

@@ 1.1 NLP, LLM, and Causality
Causal inference is an important area to uncover and leverage the causal relationships behind observations, enabling a deep understanding of the underlying mechanism and potential interventions in real-world systems. Different from most classical statistical studies, causal inference presents unique challenges due to its focus on "causation instead of correlation", which intricates a complicated integration of human knowledge (e.g., domain expertise and common sense), mathematics, and data mining. Due to the inherent proximity to the human cognitive process, causal inference has become pivotal in many high-stakes domains such as healthcare Glass et al. (2013), finance Atanasov and Black (2016), and science Imbens and Rubin (2015).
Traditional causal inference frameworks, such as structural causal model (SCM) Pearl (2009) and potential outcome framework Imbens and Rubin (2015) have systematically defined causal concepts, quantities, and measures, followed up with multiple data-driven methods to discover the underlying causal relationships Spirtes and Zhang (2016); Nogueira et al. (2022); Vowels et al. (2022) and estimate the significance of causal effects Winship and Morgan (1999); Yao et al. (2021). Despite their success, existing causal methods still fall short of matching human judgment in several key areas, such as domain knowledge, logical inference, and cultural context Kıcıman et al. (2023); Zečević et al. (2023); Jin et al. (2023a). Besides, most traditional causal inference approaches only focus on tabular data, lacking the ability to address causality in natural language. However, the demand for causal inference in natural language has persisted, offering numerous potential applications. For example, clinical text data from electronic health records (EHR) holds valuable causal information for healthcare research. Causal inference in NLP is a promising direction, offering both challenges and benefits. Advancements in large language models (LLMs) provide new opportunities to enhance traditional methods, bridging the gap between human cognition and causal inference Feder et al. (2022).
@@ 1.2 Challenges of Causal Inference in NLP
Although LLMs have shown eye-catching success, causal inference poses unique challenges for LLM capabilities. Unlike regular data, natural language is unstructured, high-dimensional, and large-scale, making traditional causal methods ineffective. Besides, causal relationships in the text are often obscure, and the complex semantics require advanced language models to uncover them. These challenges create significant hurdles for causal tasks in NLP and require new approaches, presenting transformative opportunities for advancing causal inference research.
1.3 Benefits Brought by LLMs
Despite the challenges, the increasing sophistication of LLMs has enhanced their ability in causal inference from linguistic data. LLMs bring the following key benefits to causal inference:
Domain knowledge. Traditional causal methods focus on numerical data, but domain knowledge is crucial in fields like medicine for identifying causal relationships. LLMs can extract this knowledge from large-scale text, reducing dependence on human experts for causal inference.
Common sense. LLMs can capture human common sense, which aids causal reasoning across contexts. For instance, legal cases require logic, and common sense often identifies abnormal events as causes Kıcıman et al. (2023).
Sematic concept. Natural language, with its nuances and complexity, presents challenges for identifying causal relationships. Advances in NLP and LLMs, particularly in semantic modeling, offer new opportunities for deeper causal analysis.
Explainable causal inference. LLMs can provide tools for more intuitive, natural language-based explanations of causal reasoning, making complex concepts more accessible and enhancing user interaction with causal inference results.
@@ 1.4 Contribution and Uniqueness
Contribution. This survey systematically reviews existing research on using LLMs for causal inference with main contributions including: (1) We propose a clear categorization of studies, organized by tasks (Section 2) and technologies (Section 3). (2) We present a detailed comparison of existing LLMs (Section 4) and highlight key insights, connections, and observations. (3) We provide a comprehensive summary of benchmark datasets, focusing on key aspects for further study (Table 1). (4) We identify limitations and future research directions (Section 5), offering new perspectives on underexplored areas and opportunities.
Differences from existing surveys. Several previous surveys cover related topics Liu et al. (2024); Kıcıman et al. (2023), while our survey differs as follows: (1) Main scope: We focus on "LLMs for causality," while other surveys with similar topic like Liu et al. (2024), focus more on "causality for LLMs." (2) Structure and content: Our survey uniquely organizes research around tasks, methods, datasets, and evaluation, offering a clearer and more comprehensive review. (3) Up-to-date: We include the latest advancements, providing an up-to-date perspective on current trends and progress.
@@
4.1) [ Pairwise Causal Discovery [ GPT-4 (Jin et al.; Cai et al.; Liu et al.; Gao et al.; Zečević et al.;
@@ 2.1 Causality
Structural causal model. Structural causal model (SCM) Pearl (2009) is a widely used model to describe the causal relationships inside a system. An SCM is defined with a triple $(U,V,F)$: $U$ is a set of exogenous variables, whose causes are out of the system; $V$ is a set of endogenous variables, which are determined by variables in $U\cup V$; $F=\{f_{1}(\cdot),f_{2}(\cdot),...,f_{|V|}(\cdot)\}$ is a set of functions (a.k.a. structural equations). For each $V_{i}\in V$, $V_{i}=f_{i}(pa_{i},U_{i})$, where “$pa_{i}\subseteq V\setminus{V_{i}}$” and “$U_{i}\subseteq U$" are variables that directly cause $V_{i}$. Each SCM is associated with a causal graph, which is a directed acyclic graph (DAG), where each node stands for a variable, and each arrow is a causal relationship.
Ladder of causation. The ladder of causation Pearl and Mackenzie (2018); Bareinboim et al. (2022) defines three rungs (Rung 1: Association; Rung 2: Intervention; Rung 3: Counterfactuals) to describe different levels of causation. Each higher rung indicates a more advanced level of causality. The first rung "Association" involves statistical dependencies, related to questions like "What is the correlation between taking a medicine and a disease?". Rung 2 "Intervention" moves further to allow interventions on variables, with questions like "If I take a certain medicine, will my disease be cured?". Rung 3 "Counterfactuals" relates to imagination or retrospection queries like "What if I had acted differently?", "Why?". Answering such questions requires knowledge related to the corresponding SCM. Counterfactual ranks the highest because it subsumes the first two rungs. A model that can handle counterfactual queries can also handle associational and interventional queries.
@@ 2.2 Causal Tasks and Related Rungs in Ladder of Causation
Causal inference involves various tasks. Figure 1 shows an overview of these tasks and their positions in the ladder of causation. We also show several examples of prompts corresponding to each rung. We list several main causal tasks as follows:
@@
Causal effect estimation. Causal effect estimation (a.k.a. treatment effect estimation) aims to quantify the strength of the causal influence of a particular intervention or treatment on an outcome. In different scenarios, researchers may focus on the causal effect of different granularities, such as individual treatment effect (ITE, i.e., treatment effect on a specific individual), conditional average treatment effect (CATE, i.e., average treatment effect on a certain subgroup of population), average treatment effect on the treated group (ATT), and average treatment effect (ATE, i.e., average treatment effect on the entire population). Besides, people are also interested in the direct/indirect causal effects in certain scenarios, such as natural direct effect (NDE), controlled direct effect (CDE), and natural indirect effect (NIE). Another task related to causal effect estimation is backdoor adjustment (BAJ), which aims to block all backdoor paths Pearl (2009) from the treatment to the outcome to exclude non-causal associations. Causal effect estimation tasks often span over Rung 2 and Rung 3.
Other tasks. There are many other tasks in causal inference. Among them, causal attribution (CA) refers to the process of attributing a certain outcome to certain events. Counterfactual reasoning (CR) investigates what might have happened if certain events or conditions had been different from what actually occurred. It explores hypothetical scenarios by considering alternative outcomes based on changes in “what if" circumstances. Causal explanation (CE) aims to generate explanations for an event, a prediction, or any causal reasoning process. This task often needs to answer causal questions in a specified human-understandable form or plain language. It is often in Rung 2 or 3, depending on the specific context. It is worth noting that, in many cases, different causal tasks may exhibit natural overlap in their scope. For instance, causal attribution and explanation commonly intersect with causal discovery and causal effect estimation. However, each task maintains a distinct focus.
@@
Table 1: Datasets for LLM-related causal inference, with publication year, applicable tasks (CD=causal discovery; Eff=effect estimation; CR=counterfactual reasoning; CE=causal explanation), dataset size (as these datasets are not in a consistent form, we show the size w.r.t. different units, where P=causal pairs; R=causal relations; S=samples; Q=questions), domain, generation process (R: real-world; S: synthetic), number of data sources, and citations.
@@ 3 Methodologies
Recent efforts (Kıcıman et al., 2023; Chen et al., 2024a; Gao et al., 2023) have explored leveraging LLMs for causal tasks. Unlike traditional data-driven or expert knowledge-based approaches, LLMs introduce novel methodologies, offering new perspectives for discovering and utilizing causal knowledge. Figure 2 lists LLMs developed or evaluated for causal tasks. We categorize current LLM methodologies for causal tasks as follows:
Prompting. Most existing works on causal reasoning with LLMs Chen et al. (2024a); Kıcıman et al. (2023); Long et al. (2023); Jin et al. (2023a) focus on prompting, as it is the simplest approach. This includes regular strategies (like basic prompting, In-Context Learning (ICL) Brown et al. (2020), and Chain-of-Thought (CoT) Wei et al. (2022)) and causality-specific strategies Jin et al. (2023a). For regular prompting, basic prompts (i.e., directly describing the question without any example or instruction) are most frequently used. There are also other efforts to devise more advanced prompting strategies. Among them, CaLM Chen et al. (2024a) has tested 9 prompting strategies including basic prompt, adversarial prompt Wallace et al. (2019); Perez and Ribeiro (2022), ICL, 0-shot CoT (e.g., “let’s think step by step” without any examples) Kojima et al. (2022), manual CoT (i.e., guide models with manually designed examples), and explicit function (EF) (i.e., using encouraging language in prompts) Chen et al. (2024a). Other works Kıcıman et al. (2023); Long et al. (2023); Gao et al. (2023); Ban et al. (2023) also design different prompt templates. These works show substantial improvement potential of prompt engineering in causal reasoning tasks. For example, studies Kıcıman et al. (2023); Chen et al. (2024a); Long et al. (2023) highlight that simple phrases like "you are a helpful causal assistant" can significantly boost performance. Additionally, there are causality-specific strategies, such as CausalCoT Jin et al. (2023a), which combine CoT prompting with causal inference principles Pearl and Mackenzie (2018).
@@
Combining traditional causal methods. A line of studies combines LLMs with traditional causal methods. Considering causal inference often heavily relies on numerical reasoning, an exploration in Ban et al. (2023) leverages LLMs and data-driven causal algorithms such as MINOBSx Li and Beek (2018) and CaMML O’Donnell et al. (2006). This method outperforms both original LLMs and data-driven methods, indicating a promising future for combining the language understanding capability of LLMs and the numerical reasoning skills of data-driven methods in complicated causal tasks. Jiralerspong et al. Jiralerspong et al. (2024) combine LLM with a breadth-first search (BFS) approach for full graph discovery. It considers each causal relation query as a node expansion process, and gradually constructs the causal graph by traversing it with BFS. This method significantly reduces the time complexity from $O(n^{2})$ to $O(n)$, where $n$ is the number of variables. While it does not require access to observational data, their experiments show that the performance can be further enhanced with observational statistics.
Knowledge augmentation. LLMs with augmented knowledge can often better execute tasks for which they are not well-suited to perform by themselves, particularly for causal tasks that require domain-specific knowledge. Pawlowski et al. (2023) introduced two types of knowledge augmentation: context augmentation, which provides causal graphs and ITEs in the prompt, and tool augmentation, offering API access to expert systems for causal reasoning. Tool augmentation performs more robustly across varying problem sizes, as the LLM relies on the API for reasoning instead of reasoning through the graph itself. LACR Zhang et al. (2024) applies retrieval augmented generation (RAG) to enhance the knowledge base of LLM for causal discovery, where the knowledge resources are from a large scientific corpus containing hidden insights about associational/causal relationships. Similar approaches Samarajeewa et al. (2024) employ causal graphs as external sources for causal reasoning. Knowledge augmentation is especially effective for domain-specific tasks, as seen in RC2R Yu et al. (2024) for financial risk analysis and CausalKGPT Zhou et al. (2024a) for aerospace defect analysis.